AllHome Bhd. is a hardware retail company with financial year runs from January 1 through December 31. In 2017, AllHome Bhd. sold 30000 units of its product at a price of RM450 per unit. The company reported estimated sales returns and allowances in 2017 of RM240000. AllHome actually purchased 32000 units of its product from its manufacturer in 2017 at an average cost of RM350 per unit Operating expenses (excluding depreciation) for AllHome Bhd. in 2017 were RM420000 and depreciation expense was RM100,000 AllHome had RM2000000 in total debt throughout 2017 . This debt carried an annual interest rate of 10%. Finally, All Home tax rate was 40% FE, Sem 1 2018/2019 (i) net sale:s (ii) cost of goods sold (iii) interest expense (iv) amount of taxes paid (v) net profit (vi) the net profit margin ratio and choose from the list, your interpretation of the solution Format87450000 Format50500000 Format : 600000 Format : 875000 Format : 6968000 Format : 0.07 1. gross profit per RM1 sale 2. net profit per RM1 sale 3, net profit per 200% sale 4. net profit per gross profit Format :4
Answer #1

(i) Price per Unit = RM450 and Units Sold = 30000

Gross Sales = 450 x 30000 = RM 13500000

LESS: Sales Return and Allowances = RM 240000

Net Sales = RM13260000

(ii) Cost of Goods Sold = Average Purchase Cost x Number of Units Sold = 350 x 30000 = RM 10500000

(iii) Interest Expense = Debt x Annual Interest Rate = 2000000 x 0.1 = RM 200000

(iv) Net Sales = 13260000

Less: Cost of Goods Sold = 10500000

Gross Profit = RM 2760000

Less: Operating Expenses = RM 420000

EBITDA = RM 2340000

Less: Depreciation Expense = RM 100000

EBIT = RM 2240000

Less: Interest Expense = RM 200000

EBT = RM 2040000

Tax @ 40 % = 0.4 x 2040000 = RM 936000

(v) Net Income = EBT - Taxes = 2040000 - 936000 = RM 1104000

(vi) Net Profit Margin = Net Profit per RM1 Sales = (Net Income / Net Sales) x 100 = (1104000 / 13260000) x 100 = 8.326 %